Pratt Corporation uses the weighted-average method in its process costing system. This month, the beginning inventory in the first processing department consisted of 500 units. The costs and percentage completion of these units in beginning inventory were:

Material cost: 7,100 and 75% completed

Conversion Cost: 5,700 and 25% completed

A total of 8,100 units were started and 7,500 units were transferred to the second processing department during the month. The following costs were incurred in the first processing department during the month:

Material Cost: 136,800

Conversion Cost: 322,400

The ending inventory was 80% complete with respect to materials and 75% complete with respect to conversion costs.
Note: Your answers may differ from those offered below due to rounding error. In all cases, select the answer that is the closest to the answer you computed. To reduce rounding error, carry out all computations to at least three decimal places

How many units are in ending work in process inventory in the first processing department at the end of the month?
A. 1,100
B. 900
C. 600
D. 7,600

Company for the month of May. The company uses the weighted-average method in its process costing system.
Units Labor and Overhead Percentage Complete
WIP May 1st: 10,000 40%

Units Started 64,000

Transferred Out 60,000

WIP May 31st 14,000 70%

All materials are added at the beginning of the process. Material is 100% complete.

The equivalent units for materials for the month of May were:
A. 60,000 units
B. 74,000 units
C. 64,000 units
D. 69,800 units
Answer #1

1) Ending work in process inventory = Beginning work in process+Unit started-Unit completed

= 500+8100-7500

Ending work in process inventory = 1100 Units

so answer is a) 1100

2) Equivalent unit of material = 60000+(14000*100%) = 74000

So answer is b) 74000 Units